A telescope can be used to enlarge the diameter of a laser beam and limit diffraction spreading. The laser beam is sent through the telescope in opposite the normal direction and can then be projected onto a satellite or the Moon. If this is done with the Otto Struve telescope, producing a 2.102.10 m diameter beam of 613613 nm light, what is the minimum angular spread of the beam?

minimum angular spread:

Neglecting atmospheric effects, what is the size of the spot this beam would make on the Moon, assuming a lunar distance of 3.84×1083.84×108 m?

size of spot on the Moon:
Answer #1

Angular spread , theta = 1.22 (lambda)/D = 1.22*613*10^-9 m/2.1 m = 3.56*10^-7 radian
where lambda is wavelength of beam and D is diameter

Spot length, Y = D +2y
theta = y/x
y = x(theta)
x is the diatnce between moon and earth
y is illuminated distance due to spreading
Y is total illuminated spot length

Y = 2.1 m + (2*3.84x10^8*3.56*10^-7 rad)
Y = 275.508 m